According to a study of political​ prisoners, the mean duration of imprisonment for 39 prisoners with chronic​ post-traumatic stress disorder​ (PTSD) was 31.5 months. Assuming that σ=45 ​months, determine a 95​% confidence interval for the mean duration of​ imprisonment, μ​, of all political prisoners with chronic PTSD. Interpret your answer in words
Answer #1

Confidence Interval :-
X̅ ± Z( α /2) σ / √ ( n )
Z(α/2) = Z (0.05 /2) = 1.96
31.5 ± Z (0.05/2 ) * 45/√(39)
Lower Limit = 31.5 - Z(0.05/2) 45/√(39)
Lower Limit = 17.3767
Upper Limit = 31.5 + Z(0.05/2) 45/√(39)
Upper Limit = 45.6233
95% Confidence interval is ( 17.3767 , 45.6233 )

We are 95% confident that the true population mean duration of imprisonment of all political prisoners with chronic PTSD lies within the interval ( 17.3767 , 45.6233 ).